A large wrecking ball is held in place by 2 light steel cables. If the tension Ta in the

horizontal cable is 580N,

a. What is the tension Tb in the cable that makes an angle of 40° with the vertical?

b. What is the mass of the wrecking ball?


1
Expert's answer
2021-01-21T13:01:58-0500

Let’s write the conditions of the equilibrium for the wrecking ball:


"\sum F_x=0, \sum F_y=0."

Let’s consider the forces that act on the wrecking ball in the horizontal "x"- and vertical "y"-direction:


"T_a-T_bsin\theta=0, (1)""T_bcos\theta-mg=0. (2)"

a) We can find the tension "T_b" in the cable that makes an angle of 40° with the vertical from the first equation:


"T_b=\dfrac{T_a}{sin\theta}=\dfrac{580\ N}{sin40^{\circ}}=902\ N."

b) We can find the mass of the wrecking ball from the second equation:


"T_b\cos\theta=mg,""m=\dfrac{T_bcos\theta}{g}=\dfrac{902\ N\cdot cos40^{\circ}}{9.8\ \dfrac{m}{s^2}}=70.51\ kg."

Answer:

a) "T_b=902\ N."

b) "m=70.51\ kg."